# Hardness of Range Avoidance and

### v1

@@ 1.1 Range Avoidance
Perhaps surprisingly, recent results suggested that the answer is likely no under plausible cryptographic assumptions. Ilango, Li, and Williams [ILW23] showed that Avoid is hard for deterministic algorithms assuming the existence of subexponentially secure indistinguishability obfuscation ($i\mathcal{O}$) and that ${\mathbf{NP}}\neq{\mathbf{coNP}}$. Chen and Li [CL24] extended this result and showed that Avoid is hard even for nondeterministic algorithms, under certain assumptions regarding the nondeterministic hardness of LWE (Learning with Errors) or LPN (Learning Parity with Noise). In addition to providing compelling evidence for the hardness of Avoid, these results establish a strong separation between deterministic and randomized algorithms (recall that there exists a trivial randomized algorithm for Avoid).
@@
This paper makes progress on both fronts. We show that Avoid is hard for nondeterministic algorithms under the existence of demi-bits generators with sufficient stretch [Rud97]. A formal definition of demi-bits generators is deferred to Section 2.1, and candidate constructions supporting their existence are discussed in Appendix B. For the purpose of this introduction, it suffices to keep in mind that demi-bits generators are a version of cryptographic pseudorandom generators secure against nondeterministic adversaries.
@@ 1.2 Proof Complexity Generators
    Following the idea of pseudorandomness in proof complexity, subsequent works [Kra04, Pic11, Kra11, Raz15, Kha22] studied the hardness of the Nisan–Wigderson generator ([NW94]) as a proof complexity generator in various settings. An influential conjecture of Razborov asserts that the Nisan–Wigderson generator based on any “sufficiently hard” function in ${\mathbf{NP}}\cap{\mathbf{coNP}}$ is a proof complexity generator against Extended Frege [Raz15, Conjecture 2]; that is, computational hardness can be transformed into proof complexity pseudorandomness.
@@ 1.3 Our Results
Our main result is that the existence of demi-bits generators implies that ${\mathrm{\textnormal{{Avoid}}}}\not\in{\mathbf{SearchNP}}$, i.e., Avoid is hard for nondeterministic search algorithms.
@@
If there exists a demi-bits generator $G:\{0,1\}^{n}\to\{0,1\}^{10n}$, then ${\mathrm{\textnormal{{Avoid}}}}\notin{\mathbf{SearchNP}}$.
@@
Building on this result, we show that for any fixed propositional proof system $\mathcal{P}$ closed under certain reductions, demi-bits generators for $\mathcal{P}$ imply proof complexity generators for $\mathcal{P}$. In particular, the existence of demi-bits generators secure against ${\mathbf{NP}}/_{\mathrm{poly}}$ implies the weaker version of Krajíček’s conjecture, providing strong evidence that the latter conjecture is true.
@@
Despite decades of interest, it has remained open whether $\mathsf{APC}_{1}$ and $\mathsf{PV}_{1}$ are actually distinct theories, i.e., whether $\mathrm{dwPHP}(\mathsf{PV})$ is unprovable in $\mathsf{PV}_{1}$. Recently, Ilango, Li, and Williams [ILW23] provided the first evidence separating the two: they showed that $\mathrm{dwPHP}(\mathsf{PV})$ is unprovable in $\mathsf{PV}_{1}$ under the assumptions that indistinguishability obfuscation ($i\mathcal{O}$) with JLS-security exists and that $\mathbf{coNP}$} is not infinitely often in $\mathbf{AM}$}. We remark that the same separation was also shown by Krajíček [Kra21], albeit under an assumption that is regarded as “unlikely” ($\mathbf{P}$} admits fixed-polynomial size circuits). In this work, we establish the same separation assuming the existence of demi-bits generators against ${\mathbf{AM}}/_{O(1)}$.
@@
theoremThmPVvsAPC Assuming there exists a demi-bits generator $G:\{0,1\}^{n}\to\{0,1\}^{\omega(n)}$ secure against ${\mathbf{AM}}/_{O(1)}$, the Dual Weak Pigeonhole Principle for polynomial-time functions $(\mathrm{dwPHP}(\mathsf{PV}))$ is not provable in $\mathsf{PV}$. (In particular, $\mathsf{APC}_{1}$ is a strict extension of $\mathsf{PV}_{1}$.)
The only property of $\mathsf{PV}_{1}$ used in our argument is the KPT witnessing theorem [KPT91], which states that if $\mathsf{PV}_{1}$ proves the dual weak pigeonhole principle for polynomial-time functions, then there exists a deterministic polynomial-time Student that wins the Student-Teacher game for solving Avoid in $O(1)$ rounds (see Section 2.5). Our separation result in Section 1.3 follows by showing that no such Student algorithm exists. Moreover, for any parameter $k=k(n)$, assuming the existence of demi-bits generators secure against ${\mathbf{AM}}/_{O(\log k)}$, we further rule out deterministic polynomial-time Students that wins the Student-Teacher game within $k$ rounds.
@@
Let $m=m(n)>n$ and $k=k(n)$ be parameters. If there exists a demi-bits generator $G:\{0,1\}^{n}\to\{0,1\}^{100km}$ secure against ${\mathbf{AM}}/_{O(\log k)}$, then there is no polynomial-time deterministic algorithm for Avoid on circuits with $n$ inputs and $m$ outputs using $k$ circuit-inversion oracle queries.
@@
Assuming demi-bits generators against ${\mathbf{NP}}/_{\mathrm{poly}}$, we show that our proof complexity generators are even pseudo-surjective against every proof system. (The precise definition of $k$-round pseudo-surjectivity is presented in 2.12.)
@@
Let $m=m(n)>n$ and $k=k(n)$ be parameters. If there exists a demi-bits generator $G:\{0,1\}^{n}\to\{0,1\}^{100km}$ secure against ${\mathbf{NP}}/_{\mathrm{poly}}$, then for every non-uniform propositional proof system $\mathcal{P}$, there is a non-uniform proof complexity generator $G_{\mathcal{P},k}:\{0,1\}^{n}\to\{0,1\}^{m}$ that is $k$-round pseudo-surjective against $\mathcal{P}$.
@@
This limitation is inherent to the generality of our result: we construct generators secure against all proof systems, whereas under the assumption ${\mathbf{E}}\not\subseteq{\mathbf{SIZE}}[2^{o(n)}]$, there exists a proof system that can prove circuit lower bounds (e.g., by simply hardwiring an axiom that certain $\mathbf{E}$}-complete language has exponential circuit complexity). Thus, under this standard hardness assumption, it is provably impossible to extend our results to the regime where $k\gg s$ and thereby obtain pseudo-surjectivity of the truth table generator. It remains an intriguing open question whether our approach can be refined to construct proof complexity generators of size $s$ that are $k$-round pseudo-surjective against specific systems such as Extended Frege, for some $k\gg s$. Such a result would imply that Extended Frege cannot prove any circuit lower bounds.
Finally, we comment on the strength of the adversaries required in our assumptions on demi-bits generators. In the main theorem (Theorem 1.1), a $\mathbf{SearchNP}$} algorithm for Avoid is transformed into a nondeterministic adversary that breaks the demi-bits generator. Since $\mathbf{SearchNP}$} is a uniform class, it suffices to assume that the generator is secure against uniform nondeterministic adversaries. In contrast, Theorem 1.4 requires the generator to be secure against ${\mathbf{AM}}/_{O(\log k(n))}$ adversaries. This is because the adversary invokes the Goldwasser–Sipser protocol [GS86] which is in $\mathbf{AM}$}, and needs to hardwire the index of the circuit-inversion query that succeeds with good probability. In Theorem 1.5, we require security against ${\mathbf{NP}}/_{\mathrm{poly}}$ adversaries, as our adversary needs to hardwire a “good” sequence of teacher responses in the student-teacher game, thus is highly non-uniform.
@@ 1.4 Perspective: Average-Case to Best-Case Reductions in Proof Complexity
An important topic in average-case complexity is worst-case to average-case reductions: reductions showing that if a problem $L$ is hard in the worst-case then a related problem $L^{\prime}$ is hard on average. Worst-case to average-case reductions are known for the Permanent [CPS99], Discrete Logarithm [BM84], Quadratic Residuosity [GM84], certain lattice problems [Ajt96], and more recently for problems in meta-complexity [Hir18]. On the other hand, “black-box” worst-case to average-case reductions are unlikely to exist for $\mathbf{NP}$}-complete problems [FF93, BT06b].
@@ 2.1 Demi-Bits Generators
Let $n,m$ be length parameters such that $n<m$. A function $G:\{0,1\}^{n}\to\{0,1\}^{m}$ is an $(s,\varepsilon)$-secure demi-bits generator if there is no ${\mathbf{NP}}/_{\mathrm{poly}}$ adversary $\mathsf{Adv}$ of size $s$ such that
@@
This paper requires demi-bits generators with large stretch and computable with small circuit complexity. In particular, we need the following assumptions: {restatable}[Demi-Bits Generators with Polynomial Stretch]assumptionPolyStretchDemiBits For every constant $c\geq 1$, there exists a family of demi-bits generators $\{g_{n}:\{0,1\}^{n}\to\{0,1\}^{n^{c}}\}$ secure against ${\mathbf{NP}}/_{\mathrm{poly}}$.
@@ 2.2 Arthur–Merlin Protocols
Let $\mathbf{AM}$} denote the set of languages with an Arthur–Merlin protocol. The round-collapse theorem of Babai [Bab85] implies that every language in $\mathbf{AM}$} actually has an Arthur–Merlin protocol with two rounds: $\mathsf{Verifier}$ sends the first message, $\mathsf{Prover}$ sends a proof, and $\mathsf{Verifier}$ decides whether $x\in L$. Hence, $\mathbf{AM}$} can be seen as a randomized version of $\mathbf{NP}$}; indeed, one can prove that ${\mathbf{AM}}={\mathbf{NP}}$ under circuit lower bound assumptions [KvM02, MV05, SU05, SU06].
@@
We need the following well-known $\mathbf{AM}$} protocol for proving lower bounds on the size of efficiently recognizable sets.
@@
Definition 2.3 (Breaking demi-bits generators by $\mathbf{AM}$} adversaries).
Let $m>n$. An $\mathbf{AM}$} adversary breaks a demi-bits generator $G:\{0,1\}^{n}\to\{0,1\}^{m}$ if both $\mathsf{Prover}$ and $\mathsf{Verifier}$ receives a common input $y\in\{0,1\}^{m}$, and:
@@
We also consider $\mathbf{AM}$} adversaries with advice:
Definition 2.4 (${\mathbf{AM}}/_{k(n)}$ adversaries).
An ${\mathbf{AM}}/_{k(n)}$ adversary is an Arthur–Merlin protocol where the $\mathsf{Verifier}$ is a probabilistic polynomial-time machine that additionally receives a $k(n)$-bit advice string $a_{n}$ (which may depend on the input length $n$ but not on the specific input $y$). The interaction on input $y$ proceeds as follows:
@@
  • If there exists an $\mathbf{AM}$} adversary breaking $G$, then there exists an ${\mathbf{NP}}/_{\mathrm{poly}}$ adversary breaking $G$ ([Adl78]).
  • For every constant $k\geq 2$, if there exists a $k$-round $\mathbf{AM}$} adversary breaking $G$, then there exists a (standard) two-round $\mathbf{AM}$} adversary breaking $G$ ([Bab85]).
2.3 $\mathbf{FNP}$} v.s. $\mathbf{SearchNP}$}
In this paper, we will distinguish between the two notions $\mathbf{FNP}$} and $\mathbf{SearchNP}$}.
Definition 2.6 ($\mathbf{SearchNP}$} [CL24]).
@@
The class of search problems solvable by nondeterministic polynomial-time algorithm is defined as $\mathbf{SearchNP}$}.
Definition 2.7 ($\mathbf{FNP}$} [CL24]).
The class of search problems defined by a polynomial-time relation, i.e., $R\in\mathbf{P}$ is defined as $\mathbf{FNP}$}.
While it is clear that ${\mathbf{FNP}}\subseteq{\mathbf{SearchNP}}$, the following example suggests that this inclusion is strict.
@@
If $\mathbf{P}\neq\mathbf{NP}$, then there is a total search problem in ${\mathbf{SearchNP}}\setminus{\mathbf{FNP}}$.

### v2

@@ 1.1 Range Avoidance
Perhaps surprisingly, recent results suggested that the answer is likely no under plausible cryptographic assumptions. Ilango, Li, and Williams [ILW23] showed that Avoid is hard for deterministic algorithms assuming the existence of subexponentially secure indistinguishability obfuscation ($i\mathcal{O}$) and that ${\mathsf{NP}}\neq{\mathsf{coNP}}$. Chen and Li [CL24] extended this result and showed that Avoid is hard even for nondeterministic algorithms, under certain assumptions regarding the nondeterministic hardness of LWE (Learning with Errors) or LPN (Learning Parity with Noise). In addition to providing compelling evidence for the hardness of Avoid, these results establish a strong separation between deterministic and randomized algorithms (recall that there exists a trivial randomized algorithm for Avoid).
@@
This paper makes progress on both fronts. We show that Avoid is hard for nondeterministic algorithms under the existence of demi-bits generators with sufficient stretch [^2] The stretchability of generic demi-bits generators is only partially understood. Recent work of Tzameret and Zhang [TZ24] shows that demi-bits generator with $1$-bit stretch $G:\{0,1\}^{n}\to\{0,1\}^{n+1}$ implies those with a sublinear bits of stretch $G^{\prime}:\{0,1\}^{n}\to\{0,1\}^{n+n^{c}}$ for any constant $0<c<1$. This is the first proof that generic demi-bits generators are stretchable at all, but it still falls short of the linear or polynomial stretch assumed in our hypothesis. [Rud97], e.g. generators $G:\{0,1\}^{n}\to\{0,1\}^{10n}$. A formal definition of demi-bits generators is deferred to Section 2.1, and candidate constructions supporting their existence are discussed in Appendix B. For the purpose of this introduction, it suffices to keep in mind that demi-bits generators are a version of cryptographic pseudorandom generators secure against nondeterministic adversaries.
@@ 1.2 Proof Complexity Generators
    Following the idea of pseudorandomness in proof complexity, subsequent works [Kra04, Pic11, Kra11, Raz15, Kha22] studied the hardness of the Nisan–Wigderson generator ([NW94]) as a proof complexity generator in various settings. An influential conjecture of Razborov asserts that the Nisan–Wigderson generator based on any “sufficiently hard” function in ${\mathsf{NP}}\cap{\mathsf{coNP}}$ is a proof complexity generator against Extended Frege [Raz15, Conjecture 2]; that is, computational hardness can be transformed into proof complexity pseudorandomness.
@@ 1.3 Our Results
Our main result is that the existence of demi-bits generators implies that ${\mathsf{\textnormal{{Avoid}}}}\not\in{\mathsf{SearchNP}}$, i.e., Avoid is hard for nondeterministic search algorithms.
@@
If there exists a demi-bits generator $G:\{0,1\}^{n}\to\{0,1\}^{10n}$, then ${\mathsf{\textnormal{{Avoid}}}}\notin{\mathsf{SearchNP}}$.
@@
Building on this result, we show that for any fixed propositional proof system $\mathcal{P}$ closed under certain reductions, demi-bits generators for $\mathcal{P}$ imply proof complexity generators for $\mathcal{P}$. In particular, the existence of demi-bits generators secure against ${\mathsf{NP}}/_{\mathrm{poly}}$ implies the weaker version of Krajíček’s conjecture, providing strong evidence that the latter conjecture is true.
@@
Despite decades of interest, it has remained open whether $\mathsf{APC}_{1}$ and $\mathsf{PV}_{1}$ are actually distinct theories, i.e., whether $\mathrm{dwPHP}(\mathsf{PV})$ is unprovable in $\mathsf{PV}_{1}$. Recently, Ilango, Li, and Williams [ILW23] provided the first evidence separating the two: they showed that $\mathrm{dwPHP}(\mathsf{PV})$ is unprovable in $\mathsf{PV}_{1}$ under the assumptions that indistinguishability obfuscation ($i\mathcal{O}$) with JLS-security exists and that $\mathsf{coNP}$} is not infinitely often in $\mathsf{AM}$}. We remark that the same separation was also shown by Krajíček [Kra21], albeit under an assumption that is regarded as “unlikely” ($\mathsf{P}$} admits fixed-polynomial size circuits). In this work, we establish the same separation assuming the existence of demi-bits generators against ${\mathsf{AM}}/_{O(1)}$.
@@
theoremThmPVvsAPC Assuming there exists a demi-bits generator $G:\{0,1\}^{n}\to\{0,1\}^{\omega(n)}$ secure against ${\mathsf{AM}}/_{O(1)}$, the Dual Weak Pigeonhole Principle for polynomial-time functions $(\mathrm{dwPHP}(\mathsf{PV}))$ is not provable in $\mathsf{PV}$. (In particular, $\mathsf{APC}_{1}$ is a strict extension of $\mathsf{PV}_{1}$.)
The only property of $\mathsf{PV}_{1}$ used in our argument is the KPT witnessing theorem [KPT91], which states that if $\mathsf{PV}_{1}$ proves the dual weak pigeonhole principle for polynomial-time functions, then there exists a deterministic polynomial-time Student that wins the Student-Teacher game for solving Avoid in $O(1)$ rounds (see Section 2.5). Our separation result in Section 1.3 follows by showing that no such Student algorithm exists.Moreover, for any parameter $k=k(n)$, assuming the existence of demi-bits generators secure against ${\mathsf{AM}}/_{O(\log k)}$, we further rule out deterministic polynomial-time Students that wins the Student-Teacher game within $k$ rounds.
@@
Let $m=m(n)>n$ and $k=k(n)$ be parameters. If there exists a demi-bits generator $G:\{0,1\}^{n}\to\{0,1\}^{100km}$ secure against ${\mathsf{AM}}/_{O(\log k)}$, then there is no polynomial-time deterministic algorithm for Avoid on circuits with $n$ inputs and $m$ outputs using $k$ circuit-inversion oracle queries.
@@
Assuming demi-bits generators against ${\mathsf{NP}}/_{\mathrm{poly}}$, we show that our proof complexity generators are even pseudo-surjective against every proof system. (The precise definition of $k$-round pseudo-surjectivity is presented in 2.12.)
@@
Let $m=m(n)>n$ and $k=k(n)$ be parameters. If there exists a demi-bits generator $G:\{0,1\}^{n}\to\{0,1\}^{100km}$ secure against ${\mathsf{NP}}/_{\mathrm{poly}}$, then for every non-uniform propositional proof system $\mathcal{P}$, there is a non-uniform proof complexity generator $G_{\mathcal{P},k}:\{0,1\}^{n}\to\{0,1\}^{m}$ that is $k$-round pseudo-surjective against $\mathcal{P}$.
@@
This limitation is inherent to the generality of our result: we construct generators secure against all proof systems, whereas under the assumption ${\mathsf{E}}\not\subseteq{\mathsf{SIZE}}[2^{o(n)}]$, there exists a proof system that can prove circuit lower bounds (e.g., by simply hardwiring an axiom that certain $\mathsf{E}$}-complete language has exponential circuit complexity). Thus, under this standard hardness assumption, it is provably impossible to extend our results to the regime where $k\gg s$ and thereby obtain pseudo-surjectivity of the truth table generator. It remains an intriguing open question whether our approach can be refined to construct proof complexity generators of size $s$ that are $k$-round pseudo-surjective against specific systems such as Extended Frege, for some $k\gg s$. Such a result would imply that Extended Frege cannot prove any circuit lower bounds.
Finally, we comment on the strength of the adversaries required in our assumptions on demi-bits generators. In the main theorem (Theorem 1.1), a $\mathsf{SearchNP}$} algorithm for Avoid is transformed into a nondeterministic adversary that breaks the demi-bits generator. Since $\mathsf{SearchNP}$} is a uniform class, it suffices to assume that the generator is secure against uniform nondeterministic adversaries. In contrast, Theorem 1.4 requires the generator to be secure against ${\mathsf{AM}}/_{O(\log k(n))}$ adversaries. This is because the adversary invokes the Goldwasser–Sipser protocol [GS86] which is in $\mathsf{AM}$}, and needs to hardwire the index of the circuit-inversion query that succeeds with good probability. In Theorem 1.5, we require security against ${\mathsf{NP}}/_{\mathrm{poly}}$ adversaries, as our adversary needs to hardwire a “good” sequence of teacher responses in the student-teacher game, thus is highly non-uniform.
@@ 1.4 Perspective: Average-Case to Best-Case Reductions in Proof Complexity
An important topic in average-case complexity is worst-case to average-case reductions: reductions showing that if a problem $L$ is hard in the worst-case then a related problem $L^{\prime}$ is hard on average. Worst-case to average-case reductions are known for the Permanent [CPS99], Discrete Logarithm [BM84], Quadratic Residuosity [GM84], certain lattice problems [Ajt96], and more recently for problems in meta-complexity [Hir18]. On the other hand, “black-box” worst-case to average-case reductions are unlikely to exist for $\mathsf{NP}$}-complete problems [FF93, BT06b].
@@ 2.1 Demi-Bits Generators
Let $n,m$ be length parameters such that $n<m$. A function $G:\{0,1\}^{n}\to\{0,1\}^{m}$ is an $(s,\varepsilon)$-secure demi-bits generator if there is no ${\mathsf{NP}}/_{\mathrm{poly}}$ adversary $\mathsf{Adv}$ of size $s$ such that
@@
This paper requires demi-bits generators with large stretch and computable with small circuit complexity. In particular, we need the following assumptions: {restatable}[Demi-Bits Generators with Polynomial Stretch]assumptionPolyStretchDemiBits For every constant $c\geq 1$, there exists a family of demi-bits generators $\{g_{n}:\{0,1\}^{n}\to\{0,1\}^{n^{c}}\}$ secure against ${\mathsf{NP}}/_{\mathrm{poly}}$.
@@ 2.2 Arthur–Merlin Protocols
Let $\mathsf{AM}$} denote the set of languages with an Arthur–Merlin protocol. The round-collapse theorem of Babai [Bab85] implies that every language in $\mathsf{AM}$} actually has an Arthur–Merlin protocol with two rounds: $\mathsf{Verifier}$ sends the first message, $\mathsf{Prover}$ sends a proof, and $\mathsf{Verifier}$ decides whether $x\in L$. Hence, $\mathsf{AM}$} can be seen as a randomized version of $\mathsf{NP}$}; indeed, one can prove that ${\mathsf{AM}}={\mathsf{NP}}$ under circuit lower bound assumptions [KvM02, MV05, SU05, SU06].
@@
We need the following well-known $\mathsf{AM}$} protocol for proving lower bounds on the size of efficiently recognizable sets.
@@
Definition 2.3 (Breaking demi-bits generators by $\mathsf{AM}$} adversaries).
Let $m>n$. An $\mathsf{AM}$} adversary breaks a demi-bits generator $G:\{0,1\}^{n}\to\{0,1\}^{m}$ if both $\mathsf{Prover}$ and $\mathsf{Verifier}$ receives a common input $y\in\{0,1\}^{m}$, and:
@@
We also consider $\mathsf{AM}$} adversaries with advice:
Definition 2.4 (${\mathsf{AM}}/_{k(n)}$ adversaries).
An ${\mathsf{AM}}/_{k(n)}$ adversary is an Arthur–Merlin protocol where the $\mathsf{Verifier}$ is a probabilistic polynomial-time machine that additionally receives a $k(n)$-bit advice string $a_{n}$ (which may depend on the input length $n$ but not on the specific input $y$). The interaction on input $y$ proceeds as follows:
@@
  • If there exists an $\mathsf{AM}$} adversary breaking $G$, then there exists an ${\mathsf{NP}}/_{\mathrm{poly}}$ adversary breaking $G$ ([Adl78]).
  • For every constant $k\geq 2$, if there exists a $k$-round $\mathsf{AM}$} adversary breaking $G$, then there exists a (standard) two-round $\mathsf{AM}$} adversary breaking $G$ ([Bab85]).
2.3 $\mathsf{FNP}$} v.s. $\mathsf{SearchNP}$}
In this paper, we will distinguish between the two notions $\mathsf{FNP}$} and $\mathsf{SearchNP}$}.
Definition 2.6 ($\mathsf{SearchNP}$} [CL24]).
@@
The class of search problems solvable by nondeterministic polynomial-time algorithm is defined as $\mathsf{SearchNP}$}.
Definition 2.7 ($\mathsf{FNP}$} [CL24]).
The class of search problems defined by a polynomial-time relation, i.e., $R\in\mathbf{P}$ is defined as $\mathsf{FNP}$}.
While it is clear that ${\mathsf{FNP}}\subseteq{\mathsf{SearchNP}}$, the following example suggests that this inclusion is strict.
@@
If $\mathbf{P}\neq\mathbf{NP}$, then there is a total search problem in ${\mathsf{SearchNP}}\setminus{\mathsf{FNP}}$.